OpenStudy (anonymous):

Find the volume in the first octant bounded by the coordinate planes, y=4, the plane (x/3) +(z/5) =1.

OpenStudy (anonymous):

Double integral problem

OpenStudy (anonymous):

Y end points is 0 to 4, I can solve for z to find the top what about x?

OpenStudy (dumbcow):

My guess would be since we are talking about the first octant that all values have to be positive, x goes from 0 to 3 -> z = 5-(5/3)x z bounded from 0 to 5 ->x = 3 - (3/5)z x bounded from 0 to 3

OpenStudy (anonymous):

How did you determine x?

OpenStudy (dumbcow):

i might be oversimplifying this but it seems that this forms a right triangular prism where actually you could do it without integrals V = (1/2)*5*3*4 = 30 If im mistaken i apologize

OpenStudy (anonymous):

I understand. It is an exercise: they are trying to teach us how to set up double integrals.

OpenStudy (dumbcow):

oh ok To determine x look at the linear relationship between z and x, x has to be between 0 and 3 for z to be positive

OpenStudy (anonymous):

thanks, i'll take a closer look